OpenStudy (anonymous):

I honestly do not know how to do this :( If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the second equation is substituted into the first equation. 2x - 7y = 4 3x + y = -17

OpenStudy (johnweldon1993):

So it is asking us to solve for 'y' in that second equation \[3x + y = -17\] We want to get 'y' alone on that left side of the equation...so lets subtract 3x from both sides of that equation \[y = -3x - 17\] Now you would substitute that into the FIRST equation you have....wherever you see a 'y'...

OpenStudy (johnweldon1993):

so the first equation we have is \[2x - 7y = 4 \] after substituting in the 'y' we just found we have \[2x - 7(-3x - 17) = 4\] Would be the new equation that we get

OpenStudy (johnweldon1993):

Hope that is understandable?

OpenStudy (anonymous):

YES thank you so much!

OpenStudy (johnweldon1993):

No problem!